Determine the ratio of the electric force to the gravitational force between an electron and a proton when separated by a distance r.

1
Expert's answer
2022-02-19T06:55:10-0500


Mass of electron, me"=9.11\u00d710^{-31}kg"

Mass of proton"m_p= 1.67\u00d710^{-27}kg"


Newton Gravitational constant"G= 6.67\u00d710^{-11}Nm^2kg^{-2}"



Gravitational force "F_G="

"=\frac{GM_pM_e}{r^2}=(6.67×10^{-11})\>\>\frac{1.67×10^{-27}×9.1×10^{-31}}{r^2}"


"=\frac{1.015×10^{-67}}{r^2}"


Electric force "F_e=\frac{Kq_eq_e}{r^2}=\frac{9×10^9×(1.6×10^{-19})^2}{r^2}"


"=\frac{2.304×10^{-28}}{r^2}"




"\frac{F_e}{F_q}=2.27×10^{39}"